Q.1. (x+1)= 2*log(2^x+3)-2*log(1980-2^-x)


first log has base 2 and second one has base 4. so, question is sum of roots of the  above equation


(a)3954 (b)log11(base=2)  (c)log3954(base=2)


Q.2. lim ((x^(x^x))-x^x) =? 


x tending to 0+

Grade:12

1 Answers

jitender lakhanpal
10 years ago

a.1)    option b

2 log(4) (1980 - 2^-x)    (4) means base  it can be written as 2/2 log(2) (1980 - 2^-x )  by log properties

putting the value 2^x = y finally we get the equation as

log {(y + 3 )}^2   - log {(1980 y -1)/y} = x + 1

logs are all base 2 now applying the log properties again we get

log [{y (y + 3)^2}/(1980y - 1)] = x + 1

y (y + 3)^2}/(1980y - 1) = 2 ^(x+1) = 2^x . 2

as 2^x = y we get

y^2 - 3954 y + 11 = 0

the roots will be in form of 2^x1 and 2^x2

there product   2^x1*2^x2 = 2^(x1+x2) = 11

x1+x2 = log (2) 11

 

q.2) lim x tending to 0+ x ^ x = 1

now putting this result in the limits and applying the properties of limits we get

-1 as answer

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

Thanks and Regards

Jitender

(askIITians expert)